OpenStudy (anonymous):

11m^7 4m^12 = 15^-5?

OpenStudy (anonymous):

uh, are you sure that's right?

OpenStudy (anonymous):

OH

OpenStudy (anonymous):

i messed up the question. to the -12 i meant

OpenStudy (anonymous):

What's the question? Find m?

OpenStudy (anonymous):

and it's simplify. not solve.

OpenStudy (anonymous):

oh, okay..

OpenStudy (anonymous):

11m^7*4m^-12 = 15^-5 ?

OpenStudy (anonymous):

that's what I got. so 1/15^5

OpenStudy (anonymous):

is that right question?

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

is it (11m)^7 or m^7 ?

OpenStudy (anonymous):

the first.

OpenStudy (anonymous):

I'm not solving for M but thank you.

OpenStudy (anonymous):

your first answer was what I was looking for :)

OpenStudy (anonymous):

I think you stop here: 44m^-5 = 15^-5

OpenStudy (anonymous):

(11m)^7/(4m)^12=1/(15)^5 (11)^7/(4)^12 m^5 = 1/(15)^5 how about this?

OpenStudy (anonymous):

that's what Ii got din :)

OpenStudy (anonymous):

then let it finish :) i use calculator 1,162 /m^5=1/759375